A jeweler has 2 bars of gold alloy in stock, one of 12 carats and the other of 18 carats (24 carats is pure gold, 12 carats is 12/24 pure, ect.). How many grams of each alloy must be mixed to obtain 10 grams of 14 carat gold?
I know the equation is something like x + y = 10 and 12/24x + 18/24y =14/24(10) but thats as far as I can get.

Answer by stanbon(75887)

A jeweler has 2 bars of gold alloy in stock, one of 12 carats and the other of 18 carats (24 carats is pure gold, 12 carats is 12/24 pure, ect.). How many grams of each alloy must be mixed to obtain 10 grams of 14 carat gold?
---------------------------
I know the equation is something like
x + y = 10
12/24x + 18/24y =14/24(10)
----------------------
12x + 12y = 12*10
12x + 18y = 14*10
------------------------------
subtract and solve for "y":
6y = 2*10
y = 3 1/3 grams (amt. of 18 carat gold needed)
---
Solve for "x":
x + y = 10
x = 6 2/3 grams (amt. of 12 carat gold needed)
=================================================
Cheers,
Stan H.
============